OpenStudy (anonymous):

Find the exact area of the region bounded by two concentric circles with radii 10 inches and 6 inches. a) 16pi b) 64pi c) 88pi

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

big area minus small earea

OpenStudy (anonymous):

\[\pi r^2\] big area is \[100\pi\]

OpenStudy (anonymous):

small area ?

OpenStudy (anonymous):

big area minus small area =211 right/

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think you are missing that there has to be a \(pi\) in the answer

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

\[100\pi-36\pi=?\]

OpenStudy (anonymous):

87pi

OpenStudy (anonymous):

So the answer is C?

OpenStudy (anonymous):

your arithmetic is worse than mine even !!

OpenStudy (anonymous):

Really/

OpenStudy (anonymous):

\[100-36=?\]

OpenStudy (anonymous):

Oh it's 77

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

you need a calculator

OpenStudy (anonymous):

Sorry. I know it's like 10:30 in the morning, but it takes me awhile to wake up

OpenStudy (anonymous):

yeah here too

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=100-36

OpenStudy (anonymous):

let me know when you get \(64\pi\) ...

OpenStudy (anonymous):

is this clear or no? (no is a fine answer, just asking)

OpenStudy (anonymous):

Omg I do need a calculator

OpenStudy (anonymous):

Ima go get one real quick off google

OpenStudy (anonymous):

I used to be good at this

OpenStudy (anonymous):

use the link i sent wolfram alpha is great

OpenStudy (anonymous):

not that you need a calculator to subtract \(100-36\) but still

OpenStudy (anonymous):

Oh okay nevermind then. I done already looked at it

OpenStudy (anonymous):

yeash

OpenStudy (anonymous):

I just realized that 100-30=70 and 100-36=64 cus 10-6=4

OpenStudy (anonymous):

however you like to think about it, yes

OpenStudy (anonymous):

Oh and can you help this one kid with his algebra? He's new to the site and he looks like he needs help

OpenStudy (anonymous):

sure why not?

OpenStudy (anonymous):

who is it?

OpenStudy (anonymous):

The same kid you tried to help before. He still doesn't understand what you told him and now he's calling to me for help and I don't necessarily remember how to do this stuff

OpenStudy (anonymous):

@satellite73